# Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ and Noether-Lefschetz theory of $\mathcal{A}_{g}$

### v1

@@ 1. Degree 0 invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$
We expand here explicitly the formula of Theorem 1 for the degree 00 invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to show the connections to Hilbert scheme calculations of Carlsson and Okounkov [6].
@@
[¯$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(D)$ $=$ ¯$\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
@@
Consider now the series $\big{\langle}c_{1}(\mathcal{O}/\mathcal{I})\big{\rangle}$ of [6, Corollary 3] where we take the line bundle $\mathcal{L}$ to be $\mathcal{O}$ with equivariant weight $m$ as in [6, Section 2.1.1]:
@@
The left side of (1.3) is the coefficient of $m^{1}$ of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$. By [6, Corollary 3] together with evaluations from [6, Section 2.2.2], we have
@@
by [6, Corollary 1]. The matching (1.3) then follows from a simple algebraic expansion of the $m^{1}$ coefficient of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$.
@@
[¯$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(1)$ $=$ ¯$\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
@@
which equals the $m^{1}$ coefficient of $-\frac{1}{24}\big{\langle}1\big{\rangle}$. A calculation then yields
@@ 2.2. The $n=0$ case
The $n=0$ invariant concerns the moduli space of degree 0 stable maps ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)$. After imposing the evaluation condition on the marking, the moduli space is
¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)\supset\text{ev}_{1}^{-1}(\mathsf{p}_{1})=\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}\,$
with obstruction bundle $\mathbb{E}_{g}^{\vee}\otimes\mathsf{T}$, where $\mathsf{T}$ is the tangent line on ¯$\overline{\mathcal{M}}_{1,1}$ associated to the marking.
@@
$\big{\langle}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}\big{\rangle}_{g,0}^{\pi,\circ}$ $=$ [¯$\int_{[\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}\lambda_{g}\lambda_{g-2}$
$=$ ¯$\int_{\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\lambda_{g}\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathsf{T})$
@@
Here, $\psi_{1}$ denotes the cotangent line at the marking of ¯$\overline{\mathcal{M}}_{g,1}$. In the third equality, we have denoted the dual of $c_{1}(\mathsf{T})$ by $\psi$, the cotangent line at the marking of ¯$\overline{\mathcal{M}}_{1,1}$. In the fourth equality, we have used the dilaton equation and the Hodge integral evaluated in [9, Theorem 4]. The final evaluation agrees with Theorem 2 since the Eisenstein series start with 1,
@@
$\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,0}^{\pi,\circ}$ $=$ [¯$\int_{[\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}$
$=$ ¯$\int_{\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\cdot\mathsf{e}(\mathbb{E}_{1}^{\vee}\otimes\mathsf{T})$
@@ 2.3. The $g=1$ case
For $g=1$ and $n\geq 1$, we can evaluate the integral $\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}$ by hand. There are no branch points, so the cotangent line on the domain is pulled-back from the cotangent line of ¯$\overline{\mathcal{M}}_{1,1}$. Using the latter cotangent line class, we can express $\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}$ as $\frac{1}{24}$ times a Gromov-Witten invariant of maps to a fixed elliptic target $(E,p_{1})$:
@@ 2.4. Proof of Theorem 2 for $g\geq 2$ and $n>0$
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{48}\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}.$
@@
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}\left(\mathbb{E}_{g}^{\vee}\otimes\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\right)=0\,.$
@@
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}$
@@
where $[\mathsf{pt}]\in\mathsf{CH}_{0}(\mathbb{P}^{1})$ is the point class. By Proposition 7, we obtain
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\Big{(}(-1)^{g}\lambda_{g}+(-1)^{g-1}\lambda_{g-1}\cdot 2\epsilon_{S}^{*}([\mathsf{pt}]))\Big{)}=0\,.$
@@
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g}\lambda_{g-2}$ $=$ $2\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-1}\lambda_{g-2}\cdot[\mathsf{pt}]$
@@
(2.3) [¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(E,n)]^{vir}}\tau_{1}(p_{1})\lambda_{g-1}\lambda_{g-2}=\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,,$
@@
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{24}\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,$
@@ 2.5. Proof of Proposition 7
$\mathcal{T}$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\times_{\mathbb{P}^{1}}S$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\,.$$h$$\nu$$\pi_{M}$
The target $\mathcal{T}$ is a family of elliptic curves over ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ with possible expansion over the 24 nodal fibers $R_{i}$. The only permitted expansions over $R_{i}$ are simple circuits of rational curves.
@@
$\mathcal{C}$$\mathcal{C}_{\mathsf{st}}$¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\,.$$\mathsf{st}$$\mu$$\mu_{\mathsf{st}}$
@@
Since the moduli space ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ parametrizes relative maps to the fibers of $\pi_{S}$, we have a pull-back map
@@
over ¯$\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. Since $n>0$, the map (2.4) is injective, so we obtain an exact sequence
@@
where $\mathbb{F}$ is a rank $g-1$ vector bundle on ¯$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$. We therefore have a factorization
[…]
on ¯$\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. For the second equality, the restriction of (2.5) to a fiber of $\epsilon_{S}$ is used to obtain
@@
[¯$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathsf{Tan}_{\mathbb{P}^{1}})=0$
@@ 2.6. Hodge integrals and Gromov-Witten theory for a fixed elliptic target
The integrals over ¯$\overline{\mathcal{M}}_{g,r}$ can be evaluated effectively via Hodge integral techniques.
@@
¯$\int_{\overline{\mathcal{M}}_{g,r}}\prod_{j=1}^{r}\psi_{j}^{k_{j}+1-\delta_{ij}}\cdot\lambda_{g}\lambda_{g-1}=\frac{|B_{2g}|}{2^{2g-1}(2g)!}\frac{(2g+r-3)!(2k_{i}+1)}{(2k_{1}+1)!!\ldots(2k_{r}+1)!!}\,,$
@@ 3.2. Connected/disconnected calculus
(i) Let ¯$\overline{\mathcal{M}}^{\circ}_{g}(\pi,(2,1^{n-2}))$ be the moduli space of degree $n$ stable relative maps to the fibers of the universal elliptic curve
@@
(3.2) $-\sum_{g\in\mathbb{Z}}u^{2g-3}\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\bullet}=(-i)\cdot\mathsf{trace}\left(\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(q)\right)=(-i)\cdot\mathsf{Tr}_{n}\cdot(t_{1}+t_{2}),$
after $-q=e^{iu}$.
@@
We can now state the main connected/disconnect equation which will play an essential role in the proof of Theorem 1. The proof will be presented in Section 3.4.

### v2

@@ 1. Degree 0 invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$
We expand here explicitly the formula of Theorem 1 for the degree 0 invariants of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ to show the connections to Hilbert scheme calculations of Carlsson and Okounkov [7].
@@
$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(D)$ $=$ $\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}D\cdot c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
@@
Consider now the series $\big{\langle}c_{1}(\mathcal{O}/\mathcal{I})\big{\rangle}$ of [7, Corollary 3] where we take the line bundle $\mathcal{L}$ to be $\mathcal{O}$ with equivariant weight $m$ as in [7, Section 2.1.1]:
@@
The left side of (1.3) is the coefficient of $m^{1}$ of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$. By [7, Corollary 3] together with evaluations from [7, Section 2.2.2], we have
@@
by [7, Corollary 1]. The matching (1.3) then follows from a simple algebraic expansion of the $m^{1}$ coefficient of $\langle c_{1}(\mathcal{O}/\mathcal{I})\rangle$.
@@
$\int_{[\overline{\mathcal{M}}_{1,1}(\mathsf{Hilb}^{n}(\mathbb{C}^{2}),0)]^{vir}}\text{ev}_{1}^{*}(1)$ $=$ $\int_{\overline{\mathcal{M}}_{1,1}\times\mathsf{Hilb}^{n}(\mathbb{C}^{2})}c_{\mathsf{top}}(\mathbb{E}^{*}_{1}\otimes T_{\mathsf{Hilb}^{n}(\mathbb{C}^{2})})$
@@
which equals the $m^{1}$ coefficient of $\frac{1}{24}\big{\langle}1\big{\rangle}$. A calculation then yields
@@ 2.2. The $n=0$ case
The $n=0$ invariant concerns the moduli space of degree 0 stable maps $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)$. After imposing the evaluation condition on the marking, the moduli space is
$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,0)\supset\text{ev}_{1}^{-1}(\mathsf{p}_{1})=\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}\,$
with obstruction bundle $\mathbb{E}_{g}^{\vee}\otimes\mathbb{T}$, where $\mathbb{T}$ is the tangent line on $\overline{\mathcal{M}}_{1,1}$ associated to the marking.
@@
$\big{\langle}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}\big{\rangle}_{g,0}^{\pi,\circ}$ $=$ $\int_{[\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}\lambda_{g}\lambda_{g-2}$
$=$ $\int_{\overline{\mathcal{M}}_{g,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\lambda_{g}\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathbb{T})$
@@
Here, $\psi_{1}$ denotes the cotangent line at the marking of $\overline{\mathcal{M}}_{g,1}$. In the third equality, we have denoted the dual of $c_{1}(\mathbb{T})$ by $\psi$, the cotangent line at the marking of $\overline{\mathcal{M}}_{1,1}$. In the fourth equality, we have used the dilaton equation and the Hodge integral evaluated in [10, Theorem 4]. The final evaluation agrees with Theorem 2 since the Eisenstein series start with 1,
@@
$\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,0}^{\pi,\circ}$ $=$ $\int_{[\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}]^{vir}}\psi_{1}$
$=$ $\int_{\overline{\mathcal{M}}_{1,1}\times\overline{\mathcal{M}}_{1,1}}\psi_{1}\cdot\mathsf{e}(\mathbb{E}_{1}^{\vee}\otimes\mathbb{T})$
@@ 2.3. The $g=1$ case
For $g=1$ and $n\geq 1$, we can evaluate the integral $\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}$ by hand. There are no branch points, so the cotangent line on the domain is pulled-back from the cotangent line of $\overline{\mathcal{M}}_{1,1}$. Using the latter cotangent line class, we can express $\big{\langle}\tau_{1}(\mathsf{p}_{1})\big{\rangle}_{1,n}^{\pi,\circ}$ as $\frac{1}{24}$ times a Gromov-Witten invariant of maps to a fixed elliptic target $(E,p_{1})$:
@@ 2.4. Proof of Theorem 2 for $g\geq 2$ and $n>0$
$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{48}\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}.$
@@
$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}\left(\mathbb{E}_{g}^{\vee}\otimes\epsilon_{S}^{*}(\mathsf{Tan}_{\mathbb{P}^{1}})\right)=0\,.$
@@
$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}$
@@
where $\mathsf{pt}]\in\mathsf{CH}_{0}(\mathbb{P}^{1})$ is the point class. By Proposition 7, we obtain
$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\Big{(}(-1)^{g}\lambda_{g}+(-1)^{g-1}\lambda_{g-1}\cdot 2\epsilon_{S}^{*}([\mathsf{pt}]))\Big{)}=0\,.$
@@
$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g}\lambda_{g-2}$ $=$ $2\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-1}\lambda_{g-2}\cdot[\mathsf{pt}]$
@@
(2.3) $\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(E,n)]^{vir}}\tau_{1}(p_{1})\lambda_{g-1}\lambda_{g-2}=\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,$,
@@
$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi,n)]^{vir}}\tau_{1}(\mathsf{p}_{1})\lambda_{g}\lambda_{g-2}=\frac{1}{24}\frac{|B_{2g-2}|\sigma_{2g-1}(n)}{(2g-2)!}\,$
@@ 2.5. Proof of Proposition 7
[…]
The target $\mathcal{T}$ is a family of elliptic curves over $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ with possible expansion over the 24 nodal fibers $R_{i}$. The only permitted expansions over $R_{i}$ are simple circuits of rational curves.
@@
$\mathcal{C}$$\mathcal{C}_{\mathsf{st}}$$\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)\,.$$\mathsf{st}$$\mu$$\mu_{\mathsf{st}}$
@@
Since the moduli space $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$ parametrizes relative maps to the fibers of $\pi_{S}$, we have a pull-back map
@@
over $\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. Since $n>0$, the map (2.4) is injective, so we obtain an exact sequence
@@
where $\mathbb{F}$ is a rank g1g-1 vector bundle on $\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)$. We therefore have a factorization
(2.6) $\lambda_{g}\ =\ c_{1}(\epsilon_{S}^{*}\big{(}\mathsf{Tan}_{\mathbb{P}^{1}})\big{)}\cdot c_{g-1}(\mathbb{F})\ =\ 2\epsilon_{S}^{*}(\mathsf{pt})\cdot\lambda_{g-1}\,$
on $\overline{\mathcal{M}}_{g,1}(\pi_{S},n)$. For the second equality, the restriction of (2.5) to a fiber of $\epsilon_{S}$ is used to obtain
@@
$\int_{[\overline{\mathcal{M}}^{\circ}_{g,1}(\pi_{S},n)]^{vir}}\tau_{1}(\mathsf{p_{1}})\lambda_{g-2}\cdot\mathsf{e}(\mathbb{E}_{g}^{\vee}\otimes\mathsf{Tan}_{\mathbb{P}^{1}})=0\$
@@ 2.6. Hodge integrals and Gromov-Witten theory for a fixed elliptic target
The integrals over $\overline{\mathcal{M}}_{g,r}$ can be evaluated effectively via Hodge integral techniques.
@@
$\int_{\overline{\mathcal{M}}_{g,r}}\prod_{j=1}^{r}\psi_{j}^{k_{j}+1-\delta_{ij}}\cdot\lambda_{g}\lambda_{g-1}=\frac{|B_{2g}|}{2^{2g-1}(2g)!}\frac{(2g+r-3)!(2k_{i}+1)}{(2k_{1}+1)!!\ldots(2k_{r}+1)!!}\,$,
@@ 3.2. Connected/disconnected calculus
(i) Let $\overline{\mathcal{M}}^{\circ}_{g}(\pi,(2,1^{n-2}))$ be the moduli space of degree $n$ stable relative maps to the fibers of the universal elliptic curve
@@
(3.2) $\sum_{g\in\mathbb{Z}}u^{2g-3}\langle(2,1^{n-2})\rangle_{g,n}^{E\times\mathbb{C}^{2},\bullet}=(-i)\cdot\mathsf{trace}\left(\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(q)\right)=(-i)\cdot\mathsf{Tr}_{n}\cdot(t_{1}+t_{2}),$
after $q=e^{iu}$.
@@
We can now state the main connected/disconnected equation which will play an essential role in the proof of Theorem 1. The proof will be presented in Section 3.4.